# GNSS-inertial state initialization by distance residuals

### v1

@@ IV GNSS-inertial Initialization
Figure 2: Illustration of the temporal notation for IMU and GPS measurements.
@@
and $\mathbf{R}_{k}\in SO(3)$ denotes the rotation matrix, $\mathbf{p}_{k}\in\mathbb{R}^{3}$ the position, $\bm{\omega}_{k}\in\mathbb{R}^{3}$ the angular velocity, $\mathbf{v}_{k}\in\mathbb{R}^{3}$ the linear velocity, and $\mathbf{b}^{g}\in\mathbb{R}^{3}$ the gyroscope bias. We assume a known gravity magnitude $\|\mathbf{g}\|=9.81~{}m/s^{2}$ and only optimize the gravity direction (See Sec. III-C).
@@ IV-A Cost function formulation
$\hat{\mathbf{x}}=\operatorname*{arg\hskip 1.42262ptmin\hskip 2.84526pt}_{\mathbf{x}}J(\mathbf{x})$ (9)
@@
To incorporate GNSS constraints, we consider the following GPS measurement model:
@@
Our formulation will use either global GNSS residuals when fully initialized, or relative GNSS residuals initially. The global residuals $\mathbf{r}_{\text{p}}\in\mathbb{R}^{3}$ serves as a global anchor to the fixed reference frame:
$\mathbf{r}_{\text{p}}=\mathbf{p}_{k}-\hat{\mathbf{p}}_{k}^{\text{g}},$ (14)
Relative GPS distance residuals $\mathbf{r}_{\text{d}}\in\mathbb{R}^{3}$ enforce local trajectory consistency by inter-frame motion constraints between time instants $j$ and $k$:
$\mathbf{r}_{\text{d}}=(\mathbf{p}_{j}-\mathbf{p}_{k})-(\hat{\mathbf{p}}_{j}^{\text{g}}-\hat{\mathbf{p}}_{k}^{\text{g}}),$ (15)
The global GNSS error is propagated to the relative residuals as $\bm{\Sigma}_{\text{d}}=2\bm{\Sigma}_{\text{p}}$.
@@
$\mathbf{r}_{\mathbf{g}}=\left(\frac{\mathbf{v}_{j}-\mathbf{v}_{i}}{\Delta t_{ij}}-\frac{\mathbf{R}_{i}\Delta{\mathbf{v}}_{ij}}{\Delta t_{ij}}\right)-\left(\|\mathbf{g}\|\hat{\mathbf{g}}\right),$ (16)
where $\hat{\mathbf{g}}\in{S}^{2}$ denotes the estimated gravity direction and $\|\mathbf{g}\|=9.81~{}{m/s}^{2}$ is the known gravity magnitude. Assuming the dominant uncertainty arises from inertial preintegration, the associated covariance is approximated as $\bm{\Sigma}_{\mathbf{g}}={\bm{\Sigma}_{\Delta\mathbf{v}}}/{\Delta t_{ij}^{2}}$.
@@
{dmath*}
J_i(X) = ∑_i=1^N ( ∥ rp,i ∥^2_Σp + ∥ rv,i ∥^2_Σv + ∥ r_ω,i ∥^2_Σ_ω + ∥ r_g,i ∥^2_Σ_g + ∥ rR,i ∥^2_Σϕ_ ) + ∑_j,k ∈S∥ r_d,j, k ∥^2_Σ_d
@@
where $N$ denotes the number of GPS measurements included in the optimization, and $S$ the set of GNSS pairs considered for the distance constraints. After the initialization criterion is met, the cost function is switched to the following one, that replaces the relative GNSS residuals by the global ones
{dmath*}
J(X) = ∑_i=1^N ( ∥ rp,i ∥^2_Σp + ∥ rv,i ∥^2_Σv + ∥ r_ω,i ∥^2_Σ_ω + ∥ r_g,i ∥^2_Σ_g + ∥ rR,i ∥^2_Σϕ_ + ∥ r_p,i ∥^2_Σ_p )
@@
$\mathcal{O}=\left[\begin{array}[]{cccccccc}-\mathbf{J}_{r}^{-1}(\bm{\phi})\mathbf{R}_{j}^{\top}\mathbf{R}_{i}&\mathbf{0}&\mathbf{0}&\mathbf{J}_{r}^{-1}(\bm{\phi})&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[5.81247pt] \left(\mathbf{R}_{i}^{\top}(\mathbf{v}_{j}-\mathbf{v}_{i}-\|\mathbf{g}\|\hat{\mathbf{g}}\Delta t_{ij})\right)^{\wedge}&-\mathbf{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&\mathbf{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&-\|\mathbf{g}\|\mathbf{R}_{i}^{\top}\Delta t_{ij}\\[5.81247pt] \left(\mathbf{R}_{i}^{\top}(\mathbf{p}_{j}-\mathbf{p}_{i}-\mathbf{v}_{i}\Delta t_{ij}-\tfrac{1}{2}\|\mathbf{g}\|\hat{\mathbf{g}}\Delta t_{ij}^{2})\right)^{\wedge}&-\mathbf{R}_{i}^{\top}\Delta t_{ij}&-\mathbf{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&\mathbf{R}_{i}^{\top}&\mathbf{0}&-\tfrac{1}{2}\|\mathbf{g}\|\mathbf{R}_{i}^{\top}\Delta t_{ij}^{2}\\[5.81247pt] \mathbf{0}&\mathbf{0}&-\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}\\[5.81247pt] \mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[5.81247pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}\\[5.81247pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}\\[5.81247pt] \mathbf{0}&-\tfrac{1}{\Delta t_{ij}}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\tfrac{1}{\Delta t_{ij}}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&-\|\mathbf{g}\|\mathbf{J}_{\mathbb{S}^{2}}\end{array}\right]$
@@
Figure 3: Structure of the full observability matrix $\mathcal{O}$, obtained by stacking the Jacobians of all residuals with respect to the state vector $\mathbf{x}$. The Jacobian $\mathbf{J}_{r}^{-1}(\bm{\phi})$ denotes the inverse of the right Jacobian of $\mathrm{SO}(3)$, evaluated at the rotational error $\bm{\phi}=\mathrm{Log}(\Delta\tilde{\mathbf{R}}_{ij}^{\top}\mathbf{R}_{i}^{\top}\mathbf{R}_{j})$. The last row accounts for the gravitational constraint, where the gravity direction $\hat{\mathbf{g}}\in\mathbb{S}^{2}$ is parametrized on the unit sphere and its Jacobian is denoted by $\mathbf{J}_{\mathbb{S}^{2}}$.
@@ IV-B Observability Analysis
$\mathbf{x}=\left[\mathbf{R}_{i},\mathbf{v}_{i},\mathbf{p}_{i},\;\mathbf{R}_{j},\mathbf{v}_{j},\mathbf{p}_{j},\;\mathbf{b}^{g},\hat{\mathbf{g}}\right]~{}\in\mathbb{R}^{23}$ (17)
@@
Based on the residual definitions introduced in Section IV-A, the explicit expression of the observability matrix $\mathcal{O}$ is shown in Fig. 3. Fig. 4 further illustrates its structure: each row corresponds to a measurement residual, and each column to a state variable. Nonzero entries are marked in color, with gray highlighting observable contributions and orange indicating relative-only constraints.
A symbolic evaluation indicates that the matrix $\mathcal{O}$ has a maximum rank of 23. This reduction is explained by the fact that the relative GPS residual (highlighted in orange in Fig. 4) is linearly dependent on the two absolute GPS constraints, contributing only six independent equations instead of nine.
To address this underconstrained configuration during initialization, we assume that the initial pose is known. In particular, we fix the position, velocity, and orientation at the first time step, which reduces the number of estimated variables from 23 to 14. In this reduced formulation, numerical evaluation shows that the observability matrix achieves rank 11. The remaining degrees of freedom, primarily associated with the second pose, gyroscope bias, and gravity direction, cannot be fully constrained by the available measurements at this stage.
Our approach does not rely on external priors for orientation or gravity. Instead, it leverages the intrinsic coupling between inertial preintegration and GNSS position measurements to progressively constrain the state. These simplifying assumptions are only required at the beginning of the trajectory. As the estimation window $\{t_{i}\,,t_{j}\}$ advances and new measurements are incorporated, temporal chaining and overlapping constraints from consecutive states naturally improve the system’s observability. This progressive increase in constraint information has also been analyzed in [29].
To leverage this progressive constraint accumulation, we introduce a principled criterion that delays the activation of global position residuals until sufficient motion has rendered the system adequately observable. This avoids premature incorporation of absolute GPS constraints, reducing sensitivity to early noise and improving initialization robustness.
Figure 5: Initialization with GPS measurements are shown as green squares, while IMU data points are represented as blue circles. Alignment ensures that relative motion inferred from the IMU is consistent with the global trajectory.
@@ IV-C Triggering Criterion for including Global Residuals
While the IMU provides relative motion information, GPS delivers absolute position measurements. In order to effectively integrate these two sources of information, a rigid transformation denoted as $\bm{T}_{WB}\in SE\left(3\right)$ is employed. Fig. 5 illustrates the aforementioned concept, depicting GPS measurements as green squares and IMU data points as blue circles. While GPS provides globally-referenced position information, it typically suffers from higher noise levels. In contrast, IMU measurements are available with lower short-term noise, but they are subject to significant drift over time due to biases and integration errors. To ensure a robust estimation of $\bm{T}_{WB}$, we postpone its optimization until a sufficient number of GPS measurements have been accumulated. This delay increases the system’s observability and reduces sensitivity to early-stage noise, as the integration of additional data improves excitation and allows for statistical averaging.
@@
In order to determine the triggering time instant $k^{*}$, we evaluate the observability of the extrinsic transformation $\bm{T}_{WB}$ by analyzing the singular values of an approximate Hessian matrix built from GNSS-only residuals. These residuals include one relative term (i.e., displacement between consecutive GPS positions) and two absolute anchor terms (i.e., global positions expressed in the world frame). The Hessian matrix is obtained by linearizing the aforementioned residuals with respect to the six parameters of $\bm{T}_{WB}$, yielding the following approximation:
$\mathbf{H}\approx\sum_{k=1}^{N}\mathbf{J}_{k}^{\top}\bm{\Omega}_{k}\mathbf{J}_{k},$
@@
where $\mathbf{J}_{k}\in\mathbb{R}^{9\times 6}$ is the Jacobian of the residuals at time step $k$ with respect to $\bm{T}_{WB}$, and $\bm{\Omega}_{k}=\bm{\Sigma}_{k}^{-1}\in\mathbb{S}^{9}_{+}$ is the inverse covariance matrix of the stacked GPS measurement noise. To assess the structure and conditioning of the optimization problem, we compute the singular value decomposition (SVD) of the resulting Hessian $\mathbf{H}\in\mathbb{S}^{6}_{+}$. From this decomposition, we obtain a set of singular values and define the singular value ratio as $\rho_{k}=\sigma_{\max}/\sigma_{\min}$, where $\sigma_{\min}$ and $\sigma_{\max}$ denote the smallest non-zero and the largest singular values of $\mathbf{H}$, respectively. Rather than incorporating global residuals prematurely, we defer their inclusion until the estimation problem becomes sufficiently constrained. To decide when to activate global alignment, we monitor the evolution of the relative change in the singular value ratio:
$\Delta\rho_{k}=\left|\frac{\rho_{k}-\rho_{k-1}}{\rho_{k-1}}\right|,$ (19)
@@
A large value of $\Delta\rho_{k}$ indicates that the observability of the system is still evolving, whereas consistently low values suggest that the structure of the Hessian has stabilized and the states are observable. Based on this, we introduce global GPS residuals and begin estimating the transformation $\bm{T}_{WB}$ only when the system becomes sufficiently constrained. Specifically, we activate these terms once the following condition is met:
$\Delta\rho_{k}<\Delta\rho_{\mathrm{th}}~{},$
@@
where $\Delta\rho_{\mathrm{th}}$ is a fixed threshold. The index $k^{*}$ denotes the time at which this criterion is satisfied. This activation rule complements our observability analysis by providing a concrete and data-driven signal for when global residuals can be safely introduced. In contrast to prior methods that rely on heuristic indicators or manually tuned thresholds [30, 31, 28], our approach monitors the evolution of the Hessian’s conditioning to make this decision autonomously.

### v2

@@ IV GNSS-inertial Initialization
Figure 2: Illustration of the temporal notation for IMU and GNSS measurements.
@@
and $\mathbf{R}_{k}\in SO(3)$ denotes the rotation matrix, $\mathbf{p}_{k}\in\mathbb{R}^{3}$ the position, $\bm{\omega}_{k}\in\mathbb{R}^{3}$ the angular velocity, $\mathbf{v}_{k}\in\mathbb{R}^{3}$ the linear velocity, and $\mathbf{b}^{g}\in\mathbb{R}^{3}$ the gyroscope bias. We assume a known gravity magnitude $\|\mathbf{g}\|=9.81~\mathrm{m/s^{2}}$ and only optimize the gravity direction (See Sec. III-C). The optimization is initialized with the identity quaternion, and the short initial window jointly refines attitude, gravity, and IMU biases; IMU rotational increments together with GNSS inter-epoch baselines provide sufficient yaw information for convergence without requiring an external heading prior. It is worth noting that the proposed method does not require high-rate GNSS measurements. The key requirement is the availability of inter-epoch baselines that are sufficiently informative, which depends on platform motion rather than on GNSS sampling frequency.
@@ IV-A Cost function formulation
$\hat{\mathbf{x}}=\operatorname*{arg\hskip 1.42262ptmin\hskip 2.84526pt}_{\mathbf{x}}J(\mathbf{x}).$ (8)
@@
To incorporate GNSS constraints, we consider the following measurement model:
@@
Our formulation will use either global GNSS residuals when fully initialized, and inter-epoch GNSS baseline residuals initially. Accounting for the physical offset between the IMU and the GNSS antenna, the global residuals $\mathbf{r}_{\text{p}}\in\mathbb{R}^{3}$ are defined as:
$\mathbf{r}_{\text{p}}=\mathbf{p}_{k}+\mathbf{R}_{k}\,\mathbf{l}_{B}^{\text{GNSS}}-\hat{\mathbf{p}}_{k}^{\text{g}},$ (13)
where $\mathbf{l}_{B}^{\text{GNSS}}$ is the fixed lever-arm vector from the IMU to the GNSS antenna expressed in the body frame. Inter-epoch GNSS baseline vector refers to the 3D displacement of the antenna phase center between two time instants. Inter-epoch GNSS baseline vector residuals $\mathbf{r}_{\text{d}}\in\mathbb{R}^{3}$ enforce local trajectory consistency by inter-frame motion constraints between time instants $j$ and $k$. To account for the physical offset between the IMU and the GNSS antenna, the lever arm is included in both poses, as follows
$\mathbf{r}_{\text{d}}=\left[\mathbf{p}_{j}+\mathbf{R}_{j}\,\mathbf{l}_{B}^{\text{GNSS}}-\mathbf{p}_{k}-\mathbf{R}_{k}\,\mathbf{l}_{B}^{\text{GNSS}}\right]-\left(\hat{\mathbf{p}}_{j}^{\text{g}}-\hat{\mathbf{p}}_{k}^{\text{g}}\right),$ (14)
Following [8], throughout this work, we assume that the GNSS–IMU lever arm $\mathbf{l}_{B}^{\text{GNSS}}\!\in\!\mathbb{R}^{3}$ is pre-calibrated.
The global GNSS error is propagated to the inter-epoch baseline residuals as $\bm{\Sigma}_{\text{d}}=2\bm{\Sigma}_{\text{p}}$.
@@
[…]
where $\hat{\mathbf{g}}\in{S}^{2}$ denotes the estimated gravity direction and $\|\mathbf{g}\|=9.81~\mathrm{m/s}^{2}$ is the known gravity magnitude. Assuming the dominant uncertainty arises from inertial preintegration, the associated covariance is approximated as $\bm{\Sigma}_{\mathbf{g}}={\bm{\Sigma}_{\Delta\mathbf{v}}}/{\Delta t_{ij}^{2}}$.
@@
$J_{i}(\mathcal{X})=\sum_{i=1}^{N}\left(\left\|\mathbf{r}_{\Delta\mathbf{p},i}\right\|^{2}_{\bm{\Sigma}_{\Delta\mathbf{p}}}+\left\|\mathbf{r}_{\Delta\mathbf{v},i}\right\|^{2}_{\bm{\Sigma}_{\Delta\mathbf{v}}}+\left\|\mathbf{r}_{\omega,i}\right\|^{2}_{\bm{\Sigma}_{\omega}}+\left\|\mathbf{r}_{\mathbf{g},i}\right\|^{2}_{\bm{\Sigma}_{\mathbf{g}}}+\left\|\mathbf{r}_{\Delta\mathbf{R},i}\right\|^{2}_{{\bm{\Sigma}}_{\Delta\bm{\phi}}}\right)+\sum_{j,k\in S}\left\|\mathbf{r}_{\text{d},j,k}\right\|^{2}_{\bm{\Sigma}_{\text{d}}}$
@@
where $N$ denotes the number of GNSS measurements included in the optimization, and $S$ the set of GNSS pairs considered for the distance constraints. After the initialization criterion is met, the cost function is switched to the following one, that replaces inter-epoch baseline GNSS residuals by the global GNSS ones
$J(\mathcal{X})=\sum_{i=1}^{N}\left(\left\|\mathbf{r}_{\Delta\mathbf{p},i}\right\|^{2}_{\bm{\Sigma}_{\Delta\mathbf{p}}}+\left\|\mathbf{r}_{\Delta\mathbf{v},i}\right\|^{2}_{\bm{\Sigma}_{\Delta\mathbf{v}}}+\left\|\mathbf{r}_{\omega,i}\right\|^{2}_{\bm{\Sigma}_{\omega}}+\left\|\mathbf{r}_{\mathbf{g},i}\right\|^{2}_{\bm{\Sigma}_{\mathbf{g}}}+\left\|\mathbf{r}_{\Delta\mathbf{R},i}\right\|^{2}_{{\bm{\Sigma}}_{\Delta\bm{\phi}}}+\left\|\mathbf{r}_{\text{p},i}\right\|^{2}_{\bm{\Sigma}_{\text{p}}}\right)$
@@
$\mathcal{O}=\left[\begin{array}[]{cccccccc}-\mathbf{J}_{r}^{-1}(\bm{\phi})\mathbf{R}_{j}^{\top}\mathbf{R}_{i}&\mathbf{0}&\mathbf{0}&\mathbf{J}_{r}^{-1}(\bm{\phi})&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[5.16666pt] \left(\mathbf{R}_{i}^{\top}(\mathbf{v}_{j}-\mathbf{v}_{i}-\|\mathbf{g}\|\hat{\mathbf{g}}\Delta t_{ij})\right)^{\wedge}&-\mathbf{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&\mathbf{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&-\|\mathbf{g}\|\mathbf{R}_{i}^{\top}\Delta t_{ij}\\[5.16666pt] \left(\mathbf{R}_{i}^{\top}(\mathbf{p}_{j}-\mathbf{p}_{i}-\mathbf{v}_{i}\Delta t_{ij}-\tfrac{1}{2}\|\mathbf{g}\|\hat{\mathbf{g}}\Delta t_{ij}^{2})\right)^{\wedge}&-\mathbf{R}_{i}^{\top}\Delta t_{ij}&-\mathbf{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&\mathbf{R}_{i}^{\top}&\mathbf{0}&-\tfrac{1}{2}\|\mathbf{g}\|\mathbf{R}_{i}^{\top}\Delta t_{ij}^{2}\\[5.16666pt] +\;\mathbf{R}_{i}\,[\mathbf{l}_{B}^{\text{GNSS}}]^{\wedge}&\mathbf{0}&-\mathbf{I}_{3}&-\;\mathbf{R}_{j}\,[\mathbf{l}_{B}^{\text{GNSS}}]^{\wedge}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}\\[5.16666pt] -\;\mathbf{R}_{i}\,[\mathbf{l}_{B}^{\text{GNSS}}]^{\wedge}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[5.16666pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&-\;\mathbf{R}_{j}\,[\mathbf{l}_{B}^{\text{GNSS}}]^{\wedge}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}\\[5.16666pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}\\[5.16666pt] \mathbf{0}&-\tfrac{1}{\Delta t_{ij}}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\tfrac{1}{\Delta t_{ij}}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&-\|\mathbf{g}\|\mathbf{J}_{\mathbb{S}^{2}}\end{array}\right]$
@@
Figure 3: Structure of the full observability matrix $\mathcal{O}$, obtained by stacking the Jacobians of all residuals with respect to the state vector $\mathbf{x}$. The Jacobian $\mathbf{J}_{r}^{-1}(\bm{\phi})$ denotes the inverse of the right Jacobian of $\mathrm{SO}(3)$, evaluated at the rotational error $\bm{\phi}=\mathrm{Log}(\Delta\tilde{\mathbf{R}}_{ij}^{\top}\mathbf{R}_{i}^{\top}\mathbf{R}_{j})$. The last row accounts for the gravitational constraint, where the gravity direction $\hat{\mathbf{g}}\in\mathbb{S}^{2}$ is parametrized on the unit sphere and its Jacobian is denoted by $\mathbf{J}_{\mathbb{S}^{2}}$. Accounting for the GNSS–inertial lever arm makes GNSS residuals orientation-sensitive. Absolute terms inform attitude at their node and relative baselines at both endpoints, boosting yaw/bias observability under rotation while reducing to position-only if the arm is negligible.
@@ IV-B Observability Analysis
$\mathbf{x}=\left[\mathbf{R}_{i},\mathbf{v}_{i},\mathbf{p}_{i},\;\mathbf{R}_{j},\mathbf{v}_{j},\mathbf{p}_{j},\;\mathbf{b}^{g},\hat{\mathbf{g}}\right]~\in\mathbb{R}^{23}.$ (16)
@@
$\begin{bmatrix}\frac{\partial\mathbf{r}_{\Delta\mathbf{R}}}{\partial\mathbf{x}}^{\top},\frac{\partial\mathbf{r}_{\Delta\mathbf{v}}}{\partial\mathbf{x}}^{\top},\frac{\partial\mathbf{r}_{\Delta\mathbf{p}}}{\partial\mathbf{x}}^{\top},\frac{\partial\mathbf{r}_{\text{d}}}{\partial\mathbf{x}}^{\top},\frac{\partial\mathbf{r}_{\text{p}_{i}}}{\partial\mathbf{x}}^{\top},\frac{\partial\mathbf{r}_{\text{p}_{j}}}{\partial\mathbf{x}}^{\top},\frac{\partial\mathbf{r}_{\omega}}{\partial\mathbf{x}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{g}}}{\partial\mathbf{x}}^{\top}\end{bmatrix}^{\top}$
@@
Based on the residual definitions introduced in Section IV-A, the explicit expression of the observability matrix $\mathcal{O}$ is shown in Fig. 3, while Fig. 4 illustrates its structure graphically: each row corresponds to a residual and each column to a state variable. Nonzero entries are marked in color, with gray indicating general contributions and orange highlighting relative-only constraints.
A symbolic evaluation shows that $\mathcal{O}$ has a maximum rank of 23. This is lower than the number of columns because the GNSS relative constraint is linearly dependent on the pair of absolute constraints, contributing six independent equations instead of nine. To resolve this underconstrained setup during initialization, we assume the first pose is known. Fixing position, velocity, and orientation at $t_{0}$ reduces the number of unknowns from 23 to 14. In this reduced system, $\mathcal{O}$ achieves rank 11, with the unobservable directions mainly linked to the second pose, gyroscope bias, and gravity orientation.
The actual rank of $\mathcal{O}$ depends on platform motion. In constant-velocity, straight-line trajectories, the accelerometer measures only gravity, making accelerometer bias and gravity indistinguishable. Likewise, without rotational excitation, yaw and gyroscope bias cannot be reliably separated. In both cases, the associated Jacobian columns become linearly dependent, reducing the rank of $\mathcal{O}$. This can be seen directly from our residual formulation. The velocity and position preintegration residuals ($\mathbf{r}_{\Delta v},\mathbf{r}_{\Delta p}$) contain both accelerometer bias and gravity terms ($\mathbf{R}_{i}^{\top}$ and $\|\mathbf{g}\|\mathbf{R}_{i}^{\top}\Delta t_{ij}$). With nearly constant attitude and no non-gravitational accelerations, these Jacobians collapse, causing the two states to be inseparable. Similarly, the angular velocity residual $\mathbf{r}_{\omega_{k}}$ and the rotational residual $\mathbf{r}_{\Delta R}$ constrain yaw and gyroscope bias only if the platform undergoes rotations; otherwise, their Jacobians degenerate and yaw couples with $\mathbf{b}_{g}$. As a result, certain motion patterns manifest as rank deficiencies in Fig. 3. Informative maneuvers combining accelerations and rotations are therefore essential to restore full observability [8].
Our approach does not rely on external priors for orientation or gravity. Instead, it leverages the coupling between inertial preintegration and GNSS measurements to progressively constrain the state. The simplifying assumptions at initialization are temporary: as the estimation window $\{t_{i},t_{j}\}$ advances and new measurements overlap, additional constraints accumulate and observability improves. This progressive recovery, also reported in [26], motivates our criterion that delays the activation of global GNSS residuals until the system has been sufficiently excited, thereby avoiding premature anchoring to noisy absolute positions and improving initialization robustness.
@@ IV-C Triggering Criterion for including Global Residuals
Figure 5: Initialization with GNSS measurements are shown as green squares, while IMU data are represented as blue circles. Alignment ensures that relative motion inferred from the IMU is consistent with the global trajectory.
@@
While the IMU provides relative motion information, GNSS delivers absolute position measurements. In order to effectively integrate these two sources of information, a rigid transformation denoted as $\mathbf{T}_{WB}{}\in SE\left(3\right)$ is employed. Fig. 5 illustrates the aforementioned concept, depicting GNSS measurements as green squares and IMU data points as blue circles. While GNSS provides globally-referenced position information, it typically suffers from higher noise levels. In contrast, IMU measurements are available with lower short-term noise, but they are subject to significant drift over time due to biases and integration errors. To ensure a robust estimation of $\mathbf{T}_{WB}{}$, we postpone its optimization until a sufficient number of GNSS measurements have been accumulated. This delay increases the system’s observability and reduces sensitivity to early-stage noise, as the integration of additional data improves excitation and allows for statistical averaging.
@@
In order to determine the triggering time instant $k^{*}$, we evaluate the observability of the extrinsic transformation $\mathbf{T}_{WB}{}$ by analyzing the singular values of an approximate Hessian matrix built from GNSS-only residuals. These residuals include one relative term (i.e., displacement between consecutive GNSS positions) and two absolute anchor terms (i.e., global positions expressed in the world frame). The Hessian matrix is obtained by linearizing the aforementioned residuals with respect to the six parameters of $\mathbf{T}_{WB}{}$, yielding the following approximation:
$\mathbf{H}\approx\sum_{k=1}^{N}\mathbf{J}_{k}^{\top}\bm{\Omega}_{k}\mathbf{J}_{k},$ (18)
@@
where $\mathbf{J}_{k}\in\mathbb{R}^{9\times 6}$ is the Jacobian of the residuals at time step $k$ with respect to $\mathbf{T}_{WB}{}$, and $\bm{\Omega}_{k}=\bm{\Sigma}_{k}^{-1}\in\mathbb{S}^{9}_{+}$ is the inverse covariance matrix of the stacked GNSS measurement noise. To assess the structure and conditioning of the optimization problem, we compute the singular value decomposition (SVD) of the resulting Hessian $\mathbf{H}\in\mathbb{S}^{6}_{+}$. From this decomposition, we obtain a set of singular values and define the singular value ratio as $\rho_{k}=\sigma_{\max}/\sigma_{\min}$, where $\sigma_{\min}$ and $\sigma_{\max}$ denote the smallest non-zero and the largest singular values of $\mathbf{H}$, respectively. Rather than incorporating global residuals prematurely, we defer their inclusion until the estimation problem becomes sufficiently constrained. To decide when to activate global alignment, we monitor the evolution of the relative change in the singular value ratio:
$\Delta\rho_{k}=\left|\frac{\rho_{k}-\rho_{k-1}}{\rho_{k-1}}\right|.$ (19)
@@
A large value of $\Delta\rho_{k}$ indicates that the observability of the system is still evolving, whereas consistently low values suggest that the structure of the Hessian has stabilized and the states are observable. Based on this, we introduce global GNSS residuals and begin estimating the transformation $\mathbf{T}_{WB}{}$ only when the system becomes sufficiently constrained. Specifically, we activate these terms once the following condition is met:
$\Delta\rho_{k}<\Delta\rho_{\mathrm{th}}~,$ (20)
@@
where $\Delta\rho_{\mathrm{th}}$ is a fixed threshold. The index $k^{*}$ denotes the time at which this criterion is satisfied. This activation rule complements our observability analysis by providing a concrete and data-driven signal for when global residuals can be safely introduced. In contrast to prior methods that rely on heuristic indicators or manually tuned thresholds [27, 25], our approach monitors the evolution of the Hessian’s conditioning to make this decision autonomously.